Grade 11Physical Chemistry

Find density of a 3.6M sulphuric acid that is 29% H2SO4 by mass

krutarth kamath
10 Years agoGrade 11

4 Answers

Karan Pandya
9 Years ago
Molarity of acid = 3.6;Molar mass of H2SO4 = 98gMass of H2SO4 present in 1 L (1000 mL) of acid = 3.6 x 98 g = 352.8 gLet `d` be the density of the acid.Density = Mass/Volume Mass of 1000 mL of acid = (d x 1000) g Percentage of H2SO4 in acid = 29352.9 x 100/1000 d = 29352.9/10 d = 29290 d = 352.9Therefore d = 352.9/290= 1.217 g/mL
Karan Pandya
9 Years ago
Mass of 1 liter acid =1000x density(d)29/100 x 1000d=mass of h2so4 in 1 literSo, mass of h2so4 in 1 liter= 3.6 x 98(molar mass)so, 29x10xd=3.6x98 d=1.216gm/ml
Kali
9 Years ago
As we know that,M=(%w/w ×10×d)/molar mass GivenMolarity=3.6MDensity (d) =? To findMolecular mass of H2SO4 (2×1+32+16×4)=98So 3.6=29×10×d/983.6×98/290=d So density we get is 1.22And=1.22
Riya sharma
8 Years ago
M= 10*x*D÷ molar mass where M is the molarity D is the density and x is % composition Sol.3.6= 10* 29*D÷983.6*98÷290=352.8÷290=1.216=1.22